OpenStudy (anonymous):

Find the indicated outputs for f(x)=4xsquared-4x f(0) f(-1) f(2)

OpenStudy (turingtest):

plug in the numbers f(0)=4(0)^2-4(0)=0 f(-1)=4(-1)^2-4(-1)=4+4=8 f(2)=4(2)^2-4(2)=16-8=8

OpenStudy (anonymous):

f(0)=0 f(-1)= -8 f(2)= 16 is this right?